A marble was thrown upward at an angle of elevation of 30o

. It travelled 1.2s to reach the 

highest elevation. Find its initial velocity, Vo.


1
Expert's answer
2021-06-14T08:19:23-0400
"T=\frac{v_0\sin{30}}{g}\\1.2=\frac{v_0\sin{30}}{9.8}\\v_0=23.5\frac{m}{s}"